In a negatively skewed distribution          The mean is lower than the median and...

  

  1.          In a negatively skewed distribution
  1.          The mean is lower than the median and the mode
  2.          The mean is higher than the median and the mode
  3.          The mean is between the median and mode in value
  4.          The mean is equal to the median and lower than the mode
  1.          A positively skewed distribution is caused by
  1.          A few extremely high scores
  2.          A few extremely low scores
  3.          Too many scores of intermediate value
  4.          Irregularity of measurement
  1.          What is the median for the scores: 23, 27, 29, and 13?
  1.          23        b.         25        c.         26        d.            2
Answer #1

In a negatively skewed distribution.

a. The mean is lower than the median and the mode.

--------------------------------------------------------------------------------------------------

A positively skewed distribution is caused by

a.   A few extremely high scores

----------------------------------------------------------------------------------------------------

Median is the average of middle two values when arranging in increasing order.

13,23,27,29

( 23 + 27) / 2 = 50/2 = 25

b. 25